A stock expects to pay a dividend of $3.72 per share next year. The dividend is expected to grow at 25 percent per year for three years followed by a constant dividend growth rate of 4 percent per year in perpetuity. What is the expected stock price per share 5 years from today, if the required return is 12 percent?
Answer #1

The price of a stock is the sum of present value of all its expected future dividends. The stock price in the given case can
Answer #3

What is the expected stock price per share 5 years from today, if the required return is 12 percent?
P5=D6/(r-g)=3.72*1.25^3*1.04*1.04/(12%-4%)=98.23125
Answer #2

Dividend in 5 years = dividend next year*(1+high growth period rate)^3*(1+low growth period rate)

=3.72*(1+0.25)^3*(1+0.04) = 7.56

Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
Price = 7.56 * (1+0.04) / (0.12 - 0.04)
Price = 98.28